Fermi National Laboratory


Accelerator Update

Friday September 5
The day shift began with Operations monitoring store 2988, with the Antiproton Source (Pbar) stacking, and with experts from Booster, the Main Injector (MI), and Switchyard conducting studies.

Operations terminated store 2988 at 10:31 AM.

Operations put the Tevatron (TeV) through a dry squeeze at 10:45 AM.

At 10:48 AM, the TeV suffered a sector D2 quench. Experts investigated.

Since MI and Switchyard experts were conducting septa 120 GeV studies that didn't allow stacking, a Pbar expert asked for a Pre-vault access at 1:30 PM. Operators began preparing the area for an access.

At 1:37 PM, TeV was still not on due to a sector C2 dump switch fault. A technician investigated and said it would take an hour to replace a DC breaker.

At 1:44 PM, a Booster power supply (B:MP01) tripped off. Operators reset the supply, but its current now droops when running.

The Pbar access ended at 2:46 PM. No stacking due to MI and Switchyard slow extraction studies.

Operations completed a TeV dry squeeze at 4 PM.

Operations loaded the TeV with a 36x0 store for damper studies at 4:37 PM.

Pbar resumed stacking at 4:54 PM, at a very low rate. At 5:16 PM, Pbar resumed normal stacking.

At 5:49 PM, MI vacuum valves closed. The valves opened okay, but this happened six more times over the next three hours. Operators noted that if MI wasn't stacking while the MI septa was in the beam line (the septa must be in for 120 GeV slow extraction), the vacuum looked good.

TeV studiers reported that the TeV damper system was working by 9:24 PM.

Operations pulled out the MI septa at 9:52 PM. The vacuum became stable.

Operations put the TeV through a dry squeeze at 10:28 PM.

Stacking resumed at 11:58 PM.

Saturday September 6
The midnight shift began with Pbar stacking, with Operations tuning the TeV in preparation of a squeeze.

Operations put the TeV into a shot setup at 1:30 AM.

At 3:23 AM, the TeV beam aborted due to a power glitch at the Kautz Road substation. The Duty Electrician reported that Kautz Road glitch was caused by a ComEd glitch. This tripped off many CHL cryo compressors, and the TeV RF, among many other things.

At 5:25 AM, Operations lost its computer link to MI sector 2. Operations called in experts.

CHL technicians reported at 5:28 AM that all their compressors were back on and that the transfer lines should be cold by 8 AM.

The Controls expert reestablished the computer link to MI2. He discovered a bad CAMAC crate and had it replaced by 6:51 AM.

Experts had the TeV and MI RF back on by 7:54 AM.

Operation tried to send beam the MI dump at 10:16 AM, but ran into difficulties. Operators tried various fixes with no luck. At 11:13 AM, a MI expert noted a problem with the controlling cards for the sextupole power supplies.

At 11:14 AM, Operations tried to resume Pbar stacking, but couldn't. Operators contacted an expert and then determined the problem was with the Pbar Debuncher cooling. An expert came in to make repairs.

Operations put the TeV through a squeeze at 12:38 PM.

Stacking resumed at 1:12 PM.

Operations established a 36x0 study-store at 2:48 PM. Operations terminated the beam at 5:14 PM.

Operations put the TeV into shot setup at 5:56 PM and at 8:25 PM established store 2997 with an initial luminosity of 4.1E31.

A Switchyard expert began a study period at 8:40 PM.

Operators set up Pbar to take reverse protons at 8:51 PM.

A MI expert began 2.5 MHz studies at 9:36 PM.

Pbar resumed stacking at 11:19 PM. Then an expert began a study period at 11:41 PM.

Sunday September 7
The midnight shift began with Operations monitoring store 2997, and with many studies occurring.

Pbar and Switchyard studies ended around 12:18 AM.

Operations turned off Linac at 12:19 AM in preparation for an access.

The Recycler stored some beam at 12:21 AM.

Operations locked off the Linac and Booster beam lines at 1:16 AM in preparation for the shutdown.

At 2:30 AM, a Linac RF station (LRF1) tripped off. Operators could not reset the station. An expert investigated the problem later in the morning.

MI was made ready for the shutdown around 1 PM.

Operations terminated store 2997 at 4:02 PM.

Operations made the Collider Detector Facility (CDF) and D-zero (D0) ready for access at 4:13 PM.

By 8 PM Operations had made ready all of the following areas for the shut down: Linac, Booster, MI, MI-8, MI-12A & B, TeV, Prevault, Pretarget, Transport, and Pbar rings were rad surveyed.

At 8:50 PM, a TeV expert reported completing all smart bolt measurements. This allowed cryo technicians to begin warming the TeV cryo systems at 9:41 PM.

Monday September 8
The midnight shift began with Operations continuing to prepare for the accelerator complex shutdown.

A scheduled power outage began at 7 AM. The power was restored sixteen minutes later.

Today's Plan
The plan for the day and for the next seven to ten weeks is to conduct equipment installation, maintenance, and upgrades.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/8/2003   email Fermilab